SRTC = 8 + 1⁄2q2 and therefore MC = q.



If the minimum point of the short-run ATC curve for all firms(existing and potential)is also the minimum point of the long-run average cost curve (LRAC), calculate the long-run equilibrium price, market quantity, and firm quantity. What is the long-run equilibrium number of firms in the industry? [4]


1
Expert's answer
2021-11-15T10:11:53-0500

SRTC = 8 + 1/2q2 (This is a function for each firm)


The equation can also be written as

TC = 16 + q2


In perfect competition, the long-run price equals average cost at the minimum of the average cost curve

Average cost = (16 + q2)/q = 16/q + q


Differentiate the equation to get the minimum average cost

dAC/dQ = -(16/q2) + 1 = 0

16/q2 = 1

q2 = 16

q = 4 units

Therefore, each firm produces 4 units of output in the long run.


p (price) = AC (At q = 4) = (16/q + q) = (16/4 + 4) = 8

Therefore the long-run equalibrium price is $8.


Demand function  = 1000 - 100P

qd= 1000 - (100*8) = 200

Therefore market quantity = 200 units


Number of firms = Total market quantity/ single firm quantity

= 200/4 = 50 firms


SOLUTIONS

Long-run equilibrium price = $8

Market quantity = 200 units

Firm quantity = 4 units

Number of firms = 50 firms